Simplify : (1+i)^2+(1-i)^2

Answer by Alan3354(69443)

Simplify : (1+i)^2+(1-i)^2
------
(1+i)^2 = 1^2 + 2i + i^2 = 2i
------
Find (1-i)^2 the same way, then add.